# X-SAM: From Segment Anything to Any Segmentation

### v1

@@ A.8 Further Discussions
Fig. 5: Visualization Results of Generic Segmentation. Visualized images are sampled from the COCO2017 Val set. More category names are omitted for better visualization.
Fig. 6: Visualization Results of Open-Vocabulary (OV) Segmentation. Visualized images are sampled from the ADE20K Val set. More category names are omitted for better visualization.
Fig. 7: Visualization Results of GCG Segmentation. Visualized images are sampled from the Open-PSG Val set.
Fig. 8: Visualization Results of Referring Segmentation. Visualized images are sampled from the RefCOCO Val set.
Fig. 9: Visualization Results of Reasoning Segmentation. Visualized images are sampled from the reasoning segmentation Val set.
Fig. 10: Visualization Results of Interactive Segmentation. Visualized images are sampled from the COCO-Interactive Val set. In each sample, the left image is the original image with the interactive prompts, and the right image is the visualized result of our method. From top to bottom, there are four types of visual interactive prompts: point, scribe, box, and mask. We visualize the category name of the interactive prompt for better visualization.
Fig. 11: Visualization Results of VGD Segmentation (Single Image). Visualized images are sampled from COCO-VGD Val set. In each sample, the left image is the original image with the visual grounded prompts, and the right image is the visualized result of our method. From top to bottom, there are four types of visual grounded prompts: point, scribe, box, and mask. We visualize the category name of the grounded prompt for better visualization.
Fig. 12: Visualization Results of VGD Segmentation (Cross Image). Visualized images are sampled from the COCO-VGD Val set. In each sample, the left image is a different image with the visual grounded prompts, and the right image is the visualized result of our method. From top to bottom, there are four types of visual grounded prompts: point, scribe, box, and mask. We visualize the category name of the grounded prompt for better visualization.

### v2

@@ Appendix A Technical Appendices and Supplementary Material
Figure 6: Visualization Results of Generic Segmentation. Visualized images are sampled from the COCO2017 Val set. More category names are omitted for better visualization.
Figure 7: Visualization Results of Open-Vocabulary (OV) Segmentation. Visualized images are sampled from the ADE20K Val set. More category names are omitted for better visualization.
Figure 8: Visualization Results of GCG Segmentation. Visualized images are sampled from the Open-PSG Val set.
Figure 9: Visualization Results of Referring Segmentation. Visualized images are sampled from the RefCOCO Val set.
Figure 10: Visualization Results of Reasoning Segmentation. Visualized images are sampled from the reasoning segmentation Val set.
Figure 11: Visualization Results of Interactive Segmentation. Visualized images are sampled from the COCO-Interactive Val set. In each sample, the left image is the original image with the interactive prompts, and the right image is the visualized result of our method. From top to bottom, there are four types of visual interactive prompts: point, scribe, box, and mask. We visualize the category name of the interactive prompt for better visualization.
Figure 12: Visualization Results of VGD Segmentation (Single Image). Visualized images are sampled from COCO-VGD Val set. In each sample, the left image is the original image with the visual grounded prompts, and the right image is the visualized result of our method. From top to bottom, there are four types of visual grounded prompts: point, scribe, box, and mask. We visualize the category name of the grounded prompt for better visualization.
Figure 13: Visualization Results of VGD Segmentation (Cross Image). Visualized images are sampled from the COCO-VGD Val set. In each sample, the left image is a different image with the visual grounded prompts, and the right image is the visualized result of our method. From top to bottom, there are four types of visual grounded prompts: point, scribe, box, and mask. We visualize the category name of the grounded prompt for better visualization.